# Modelling Quantum Transduction for Multipartite Entanglement Distribution

### v1

@@ Abstract
Superconducting and photonic technologies are envisioned to play a key role in the Quantum Internet. However the hybridization of these technologies requires functional quantum transducers for converting superconducting qubits into “flying” qubits able to propagate through the network (and vice-versa). In this paper, quantum transduction is theoretically investigated for a key functionality of the Quantum Internet, namely, multipartite entanglement distribution. Different communication models for quantum transduction are provided, in order to make the entanglement distribution possible. The proposed models departs from the large heterogeneity of hardware solutions available in literature, abstracting from the particulars of the specific solutions with a communication engineering perspective. Then, a performance analysis of the proposed models is conducted through key communication metrics, such as quantum capacity and entanglement generation probability. The analysis reveals that – although the considered communication metrics depend on transduction hardware parameters for all the proposed models – the particulars of the considered transduction paradigm play a relevant role in the overall entanglement distribution performance.
@@ I Introduction
Multipartite entanglement has recently recognized as a crucial resource for enabling astonishing functionalities in the Quantum Internet [2, 3, 4]. Thus, multipartite entanglement distribution is a key research area from a communication and network engineering perspective. Nevertheless, what is missing is a consolidated literature that treats multipartite entanglement distribution with an outlook on quantum transduction, which represent a crucial challenge for the physical implementation of the Quantum Internet. In fact, the final stage of the Quantum Internet is envisioned as the hybridization of different quantum technologies [5] – such as superconducting and photonic technologies – aiming at exploiting the complementary features of each technology and combining their different strengths to achieve superior performance and reliability.
@@
Indeed, superconducting technology is recognized as a very promising quantum computing platform due to its capabilities to realize fast gates and due to its high scalability [6]. However, the superconducting technology requires cryogenic temperatures, which in turn challenge the development of large-scale quantum networks. On the other hand, photonic technology is worldwide recognized as the most suitable technology for realizing the so-called flying qubits, i.e., optical photons acting as quantum carriers, which travel along communication channels for fulfilling quantum communication needs. Optical photons weakly interact with the environment (thus, less subjected to decoherence), they can be easily controlled with standard optical components as well as they are characterized by high-rate low-loss transmissions [7, 5].
@@
However, flying qubits working at optical frequencies (typically about hundred of THz) cannot directly interact with superconducting qubits that, conversely, work at microwave frequencies (GHz). Hence, for integrating these two technologies in a quantum network, a quantum transducer is needed for converting a superconducting qubit within a network node into a flying qubit [8, 9, 10, 11, 12] and vice-versa. Yet, transduction between microwave and optical domain still represents an open problem, due to the huge frequency gap – about five orders of magnitude – between microwave and optical photons.
@@
In this paper, the transduction process is investigated in order to make the multipartite entanglement distribution possible. To this aim, we analyze the capabilities of different paradigms underlying quantum transduction, namely direct conversion and intrinsic entanglement generation. Based on these capabilities, we propose different communication models for multipartite entanglement distribution, by accounting for the main transducer-hardware parameters. The proposed models can be applied to any type of multipartite entangled state, and we theoretically compare the different models through key communication metrics, such as quantum capacity and entanglement generation probability.
The conducted analysis reveals that the aforementioned communication metrics are functions of the main transducer-hardware parameters, which hugely influence the performance of the entanglement distribution. But the analysis shows also that the particulars of the adopted transducer paradigm play a relevant role in the overall entanglement distribution performance, with a paradigms being able somehow to overcome the stringent hardware constraints imposed by other paradigms. Accordingly, we identify some crucial trade-offs in this comparison, given the current state-of-the-art technology.
@@
The remaining part of the manuscript is organised as follows. In Section II, we describe the problem statement by presenting two main communication paradigms for multipartite entanglement distribution. In Section III, we develop and analyze the different communication models for multipartite distribution, by accounting for the different transducer paradigms. Finally, in Section IV we conclude the paper.
@@
Figure 1: Direct Multipartite entanglement Distribution (DMT) for a $3$-qubit GHZ state. The multipartite state is generated locally at the superconducting orchestrator, and it must be distributed to the superconducting nodes representing the three clients via optical quantum channels. Ebits at microwave and optical frequencies are depicted in blue and red, respectively. The Quantum Transducers (QTs) at the orchestrator realize an up-conversion of the ebits of the GHZ state, by converting them from microwave to optical frequencies. After being distributed through optical channels, the ebits are converted again into microwave frequencies with a down-conversion process implemented by the QTs at the clients.
@@ II Problem statement: Direct vs Teleported Multipartite Entanglement Distribution
(a) Network state before EPR distribution: the orchestrator locally generates the multipartite state as well as additional microwave EPRs – one for each ebit of the multipartite state that must be distributed to the clients – that are distributed through up- and down-conversion processes, while no conversion are required for the multipartite state.
(b) Network state after EPR distribution: EPR pairs have been distributed (ideally) so that one microwave ebit is at the orchestrator and the other microwave ebit is at each client. By consuming the EPR pairs during the teleportation processes, each microwave ebit of the multipartite state – a $3$-qubit GHZ state in this case – is teleported at the corresponding client
Figure 2: TMD: Teleported Multipartite entanglement Distribution for the same multipartite state considered in Fig. 1. First, the orchestrator generates and share three EPR pairs with the clients, which represents the communication resource utilized for teleporting the multipartite state to the clients. Then, by performing local operations – i.e., Bell State Measurements (BSMs) – between the ebits of the EPR pairs at the orchestrator and the ebits of the GHZ state, followed by classical communications [13], the $3$-qubit GHZ is shared between 3 clients.
@@
Indeed, the generation of multipartite entanglement requires sophisticated and resource-intensive setups, often involving complex experimental apparatuses and precise control mechanisms. This makes pragmatic to assume a specialized super-node, in the following referred to as orchestrator, responsible for the entanglement generation and distribution [14, 15, 16]. The orchestrator is connected via quantum channels to network nodes with lower capabilities of satisfying entanglement technological and hardware requirements, referred to as clients. Accordingly, to eventually distribute a multipartite entangled state among the clients, the orchestrator first locally generates the multipartite entanglement state. Then, the entangled qubits – ebits in the following – of the multipartite state are distributed to the clients according to some distribution strategy [4].
It is a matter of fact that distributing multipartite entangled states among remote superconducting quantum nodes constraints the adopted distribution strategy, which must be able to account for the huge frequency gap between intra- and inter-nodes frequencies – namely, between microwave and optical frequencies – by resorting to quantum transduction. In this context, we can distinguish two main categories for multipartite entanglement distribution strategies:
@@
In DMD, as suggested by the name, the ebits are directly converted from microwave to optical frequencies and vice-versa, in order to be distributed to the clients, as shown in Fig. 1. Conversely, in TMD, the ebits are teleported to the clients, by exploiting additional EPR pairs that have been generated and shared between orchestrator and the clients, as shown in Fig. 2.
It is worthwhile to mention that, in the figures, we considered a GHZ [^1] A Greenberger–Horne–Zeilinger (GHZ) state – formally, |$\ket{GHZ}=\frac{1}{\sqrt{2}}\left(\ket{0}^{\otimes n}+\ket{1}^{\otimes n}\right)$ for a $n$-qubit state [17] – is a maximally entangled state, characterised by maximally connectivity. Indeed, a state is maximally connected if, for any two qubits, there exists a sequence of single-qubit measurements on the remaining qubits that, when performed, guarantee that the two qubits end up in a maximally entangled state [18]. On the other hand, GHZ states exhibit minimum persistency equal to $1$. Indeed, the persistency of a multipartite entangled state is the minimum number of qubits that need to be measured to guarantee that the resulting state is unentangled [18, 3]. state only for the sake of exemplification. Obviously, our modelling is not limited to this type of multipartite entanglement class, but it can be rather applied to any other class.
@@ II-A DMD: Direct Multipartite entanglement Distribution
Each ebit of the multipartite state can be directly distributed from the orchestrator to each client. This task requires to perform two different frequency conversions for each ebit of the multipartite entangled state, so that it can be mapped from superconducting qubit to flying qubit and viceversa:
@@
More into details, each ebit of the multipartite entangled state, locally generated at the orchestrator, must be distributed to remote client nodes. Unfortunately, superconducting qubits operate at microwave frequencies, which natively propagate to very limited distances, in the order of few meters [19]). Hence, interconnecting quantum nodes to distances of practical interest requires to transduce microwave frequencies into higher frequencies, such as the optical ones that allows to leverage the low-loss provided by optical fiber links. In such a way, the up-converted ebit can propagate via optical carriers to the client nodes, where it must be down-converted to microwave frequencies once again, so that it can be processed by the superconducting client.
@@
It is important to highlight that both up- and down-conversion processes are not deterministic. As a matter of fact, there exists a non-zero probability that either or both the conversions fail [20, 21, 22], with failure-probability values strictly depending on the particulars of the hardware used for implementing the microwave-optical transduction. As instance, electro-optical transduction [^2] Electro-optical conversion exhibits several attractive features from a communication perspective with respect to other platforms – such as transduction based on atomic-ensable or opto-electro-mechanics – ranging from being mechanically and thermally stable through broadband to (potentially) low-noise [9]. achieves values for the successful transduction probability – also referred to as conversion efficiency in Sec.III-A – in the order of $10^{-2}$ [10, 23] for both up- and down-conversion. And despite extensive research efforts in the realization of quantum transducers based on different platforms, obtaining high efficiency is still an open and crucial challenge [24].
Accordingly, a successful DMD requires to preserve the multipartite state, originally generated at the orchestrator, during the whole distribution process. This implies that each ebit must be preserved during both up- and down-conversion processes (as well as during the carrier propagation process through the optical link) for each client. Clearly, whenever all the aforementioned processes succeed, the original multipartite state is successfully distributed among the clients. Conversely, whenever any of the mentioned processes should fails for any of the clients, then the distribution of the entire multipartite state could be compromised. The reason for the aforementioned statement is that entanglement among the remaining ebits may or may not survive to the failure, depending on the persistence property – defined in Footnote 1 – of the specific class of multipartite entangled state to be distributed [18]. As an example, the direct distribution of GHZ-like states, which are characterized by the lowest persistence, requires all the ebits encoding the GHZ state to be successfully distributed to the clients in a single distribution attempt [4, 14, 25]. Hence, even the loss of a single ebit of the original GHZ state – during up-/down-conversion or during the transmission through the optical channel – results in the disruption of the whole multipartite state.
@@ II-B TMD: Teleported Multipartite entanglement Distribution
Differently from DMD, TMD exploits the preliminary distribution of EPR pairs to the clients for eventually distributing the multipartite entangled state via quantum teleportation protocol. Specifically, once the EPR pairs [^3] It is worthwhile to note that the generation of the EPR pairs can happen either sequentially or in parallel, depending on the characteristics of the underlying quantum technology. have been generated, the orchestrator retains one ebit of each EPR pair while distributing the other ebit to the corresponding client. Subsequently, once the EPR ebits have been successfully received by the clients, the multipartite entangled state is teleported at the clients by performing local operations and classical communications [5], as exemplified in Fig. 2.
Accordingly, TMD shifts the impact of noisy quantum transduction and noisy optical-ebit carrier propagation from multipartite ebits to EPR ebits. In fact,quantum transduction in TMD acts on EPR pairs only, and no up- or down-conversion of the multipartite ebits is required. As a consequence, this shifting makes TMD viable for all the classes of multipartite entanglement, regardless of their persistence properties. Moreover, TMD strategy guarantees more resilience to noise and better protection against memory decoherence [4, 26]. Indeed, any loss or noise would affect the EPR pair only, which – differently from the multipartite entanglement state – can be (more) easily regenerated.

### v2

@@ Abstract
Superconducting and photonic technologies are envisioned to play a key role in the Quantum Internet. However the hybridization of these technologies requires functional quantum transducers for converting superconducting qubits, exploited in quantum computation, into “flying” qubits, able to propagate through the network (and vice-versa). In this paper, quantum transduction is theoretically investigated for a key functionality of the Quantum Internet, namely, multipartite entanglement distribution. Different communication models for quantum transduction are provided, in order to make the entanglement distribution possible. The proposed models departs from the large heterogeneity of hardware solutions available in literature, abstracting from the particulars of the specific solutions with a communication engineering perspective. Then, a performance analysis of the proposed models is conducted through key communication metrics, such as quantum capacity and entanglement generation probability. The analysis reveals that – although the considered communication metrics depend on transduction hardware parameters for all the proposed models – the particulars of the considered transduction paradigm play a relevant role in the overall entanglement distribution performance.
@@ I Introduction
Multipartite entanglement has recently recognized as a crucial resource for enabling astonishing functionalities in the Quantum Internet [2, 3, 4]. Thus, multipartite entanglement distribution is a key research area from a communication and network engineering perspective. Nevertheless, what is missing is a consolidated literature that treats multipartite entanglement distribution with an outlook on quantum transduction, a crucial challenge towards the vision of the Quantum Internet. Indeed, the final stage of the Quantum Internet is envisioned as the hybridization of different quantum technologies [5] – such as superconducting and photonic technologies – aiming at exploiting the complementary features of each technology and combining their different strengths to achieve superior performance and reliability.
@@
More into details, on one hand, superconducting technology is recognized as a very promising quantum computing platform because of its capabilities to realize fast gates and its high scalability [6]. However, the superconducting technology requires cryogenic temperatures, which in turn challenge the development of large-scale quantum networks. On the other hand, photonic technology is worldwide recognized as the most suitable technology for realizing the so-called flying qubits, i.e., optical photons acting as quantum carriers, which travel along communication channels for fulfilling quantum communication needs. In fact, optical photons weakly interact with the environment (thus, less subjected to decoherence), they can be easily controlled with standard optical components and they are characterized by high-rate low-loss transmissions [7, 5].
@@
However, flying qubits working at optical frequencies (typically about hundred of THz) cannot directly interact with superconducting qubits that, conversely, work at microwave frequencies (GHz). Hence, for integrating these two technologies in a quantum network, a quantum transducer is needed for converting a superconducting qubit within a network node into a flying qubit that travels through optical channels [8, 9, 10, 11, 12] and vice-versa. Yet, transduction between microwave and optical domain still represents an open problem, due to the huge frequency gap – about five orders of magnitude – between microwave and optical photons.
@@
In this paper, the transduction process is investigated in order to make the multipartite entanglement distribution possible. To this aim, we analyze the capabilities of two different paradigms underlying quantum transduction, namely direct conversion and intrinsic entanglement generation. Stemming from these transduction paradigms, we propose different communication models for multipartite entanglement distribution, that can be applied to any class of multipartite entangled state. We theoretically compare the proposed models through key communication metrics, such as quantum capacity and entanglement generation probability. The conducted analysis reveals that the aforementioned communication metrics are functions of the main transducer-hardware parameters, which hugely influence the performance of the entanglement distribution. Accordingly, we identify some crucial trade-offs in this comparison, given the current state-of-the-art technology. The identified trade-offs are expressed in terms of achievable quantum capacity and required cooperativity (defined in Sec. III-A), as deeply discussed in Sec. IV and summarized in Tab. III in that section.
@@
The remaining part of the manuscript is organised as follows. In Section II, we describe the problem statement by presenting two main communication paradigms for multipartite entanglement distribution. In Section III, we develop and analyze the different communication models for multipartite distribution, by accounting for the different transducer paradigms. In Section IV, we compare the proposed models and finally, in Section V we conclude the paper.
@@
Figure 1: Direct Multipartite entanglement Distribution (DMD) for a $3$-qubit GHZ state. The multipartite state is generated locally at the superconducting orchestrator, and it must be distributed to the superconducting nodes representing the three clients via optical quantum channels. Ebits at microwave and optical frequencies are depicted in blue and red, respectively. The Quantum Transducers (QTs) at the orchestrator realize an up-conversion of the ebits of the GHZ state, by converting them from microwave to optical frequencies. After being distributed through optical channels, the ebits are converted again into microwave frequencies with a down-conversion process implemented by the QTs at the clients.
@@ II Problem statement: Direct vs Teleported Multipartite Entanglement Distribution
(a) Network state before EPR pairs distribution: the orchestrator locally generates the multipartite state as well as additional microwave EPR pairs – one for each ebit of the multipartite state that must be distributed to the clients – that are distributed through up- and down-conversion processes, while no conversion are required for the multipartite state.
(b) Network state after EPR pairs distribution: EPR pairs have been distributed (ideally) so that one microwave ebit is at the orchestrator and the other microwave ebit is at each client. By consuming the EPR pairs during the teleportation processes, each microwave ebit of the multipartite state – a $3$-qubit GHZ state in this case – is teleported at the corresponding client
Figure 2: TMD: Teleported Multipartite entanglement Distribution for the same multipartite state considered in Fig. 1. First, the orchestrator generates and share three EPR pairs with the clients, which represent the communication resource utilized for teleporting the multipartite state to the clients. Then, by performing local operations – i.e., Bell State Measurements (BSMs) – between the ebits of the EPR pairs at the orchestrator and the ebits of the GHZ state, followed by classical communications [13], the $3$-qubit GHZ is shared between 3 clients.
@@
Indeed, the generation of multipartite entanglement requires sophisticated and resource-intensive setups, often involving complex experimental apparatuses and precise control mechanisms. This makes pragmatic to assume a specialized super-node, in the following referred to as orchestrator, responsible for the entanglement generation and distribution [14, 15, 16]. The orchestrator is connected via quantum channels to network nodes with lower capabilities of satisfying entanglement technological and hardware requirements, referred to as clients. Accordingly, to eventually distribute a multipartite entangled state among the clients, the orchestrator first locally generates the multipartite entanglement state. Then, the entangled qubits – ebits in the following – of the multipartite state are distributed to the clients according to selected distribution strategies [4].
In the considered scenario, the orchestrator is a superconducting node, performing quantum computations on multipartite entangled qubits. The rational for this design setup choice lies in its killer application, namely the distributed quantum computation [17] in a quantum data center (also known in literature as quantum server farm) [18]. Indeed, this constitutes a very hot research topic, which is gathering a huge attention by technology giants such as Amazon [19], IBM [20] and Cisco [21, 22, 23]. This attention is not surprising, since, in a short-term time-horizon, interconnecting different quantum processors with a quantum local-area network (Q-LAN) represents the only way (with current technology-readiness level) to scale the number of qubits beyond the thousands that are necessary to achieve fault-tolerant quantum computing [16, 17]. Recently we showed in [16, 24, 25, 26] that, by leveraging the local operations at the orchestrator, it is possible to boost the network performance in terms of overhead and reliability with respect to the scenario without a supernode.
In this context, distributing multipartite entangled states among remote superconducting quantum nodes constraints the adopted distribution strategy. Indeed, the chosen strategy must take into account the huge frequency gap between intra- and inter-nodes frequencies – namely, between microwave and optical frequencies – by resorting to quantum transduction. And, we can distinguish two main categories for multipartite entanglement distribution strategies:
@@
In DMD, as suggested by the name, the ebits of the multipartite state are directly converted via quantum transducers (QTs) from microwave to optical frequencies and vice-versa, in order to be distributed to the clients, as shown in Fig. 1. Conversely, in TMD, the ebits of the multipartite state are never transducer, but they are teleported to the clients. The quantum teleportation of the multipartite state exploits EPR pairs that have been proactively generated and shared between orchestrator and the clients [27], as shown in Fig. 2. Thus in TMD, the ebits of the EPR pairs are the only ones subjected to the transduction process. The main differences between DMD and TMD are summarized in Tab. I.
It is worthwhile to mention that, in the figures, we considered a $3$-qubit GHZ [^1] A Greenberger–Horne–Zeilinger (GHZ) state – formally, |$\ket{GHZ}=\frac{1}{\sqrt{2}}\left(\ket{0}^{\otimes n}+\ket{1}^{\otimes n}\right)$ for a $n$-qubit state [28] – is a maximally entangled state, characterised by maximally connectivity. Indeed, a state is maximally connected if, for any two qubits, there exists a sequence of single-qubit measurements on the remaining qubits that, when performed, guarantee that the two qubits end up in a maximally entangled state [29]. On the other hand, GHZ states exhibit minimum persistency equal to $1$. Indeed, the persistency of a multipartite entangled state is the minimum number of qubits that need to be measured to guarantee that the resulting state is unentangled [29, 3]. state to be distributed to the clients. This choice is made only for the sake of pictorial representation, since our modelling is not limited to this class of multipartite entanglement or to the number of clients. Indeed, the proposed strategies hold for a generic $n$-qubit multipartite state to be distributed to $n$ clients [^2] To be more precise, the proposed analysis and strategies work also for the general case where the number of client is $m\leq n$. This results into a multipartite entangled state shared between the orchestrator and clients, i.e., the orchestrator retains some qubit at its side as in [16, 24, 25, 26]. However, in the following, for the sake of pictorial representation, we consider the scenario where the number of ebit of the multipartite state coincides with the number of clients..
TABLE I: DMD versus TMD strategy for multipartite entanglement distribution.
@@ II-A DMD: Direct Multipartite entanglement Distribution
In DMD each ebit of the multipartite state can be directly distributed from the orchestrator to each client. This task requires QTs that perform two different frequency conversions for each ebit of the multipartite entangled state, so that it can be mapped from superconducting qubit to flying qubit and viceversa:
@@
More into details, each ebit of the multipartite entangled state, locally generated at the orchestrator, must be distributed to remote client nodes. Unfortunately, superconducting qubits operate at microwave frequencies, which natively propagate to very limited distances, in the order of few meters [30] . Hence, interconnecting quantum nodes to distances of practical interest requires to transduce microwave frequencies into higher frequencies, such as the optical ones that allows to leverage the low-loss provided by optical fiber links. In such a way, the up-converted ebit can propagate via optical carriers to the client nodes, where it must be down-converted to microwave frequencies once again, so that it can be processed by the superconducting client.
@@
It is important to highlight that both up- and down-conversion processes are not deterministic. As a matter of fact, there exists a non-zero probability that either or both the conversions fail [31, 32, 33], with failure-probability values strictly depending on the particulars of the hardware used for implementing the microwave-optical transduction. As instance, electro-optical transduction [^3] Electro-optical conversion exhibits several attractive features from a communication perspective with respect to other platforms – such as transduction based on atomic-ensable or opto-electro-mechanics – ranging from being mechanically and thermally stable through broadband to (potentially) low-noise [9]. achieves values for the successful transduction probability – also referred to as conversion efficiency in Sec.III-A – in the order of $10\%$ [34] for both up- and down-conversion. And despite extensive research efforts in the realization of quantum transducers based on different platforms, obtaining high efficiency is still an open and crucial challenge [35].
Accordingly, a successful DMD requires to preserve the multipartite state, originally generated at the orchestrator, during the whole distribution process. This implies that each ebit must be preserved during both up- and down-conversion processes (as well as during the carrier propagation process through the optical link) for each client. Clearly, whenever all the aforementioned processes succeed, the original multipartite state is successfully distributed among the clients. Conversely, whenever any of the mentioned processes should fails for any of the clients, then the distribution of the entire multipartite state could be compromised. This vulnerability should be avoided, as multipartite states are much more complex to generate and control with respect to bipartite entangled states. The reason for the aforementioned statement is that entanglement among the remaining ebits may or may not survive to the failure, depending on the persistence property – defined in Footnote 1 – of the specific class of multipartite entangled state to be distributed [29]. As an example, the direct distribution of GHZ-like states, which are characterized by the lowest persistence, requires all the ebits encoding the GHZ state to be successfully distributed to the clients in a single distribution attempt [4, 14, 36]. Hence, even the loss of a single ebit of the original GHZ state – during up-/down-conversion or during the transmission through the optical channel – results in the disruption of the whole multipartite state.
@@ II-B TMD: Teleported Multipartite entanglement Distribution
Differently from DMD, TMD exploits the preliminary distribution of EPR pairs to the clients for eventually distributing the multipartite entangled state via quantum teleportation protocol. Specifically, once the EPR pairs [^4] It is worthwhile to note that the generation of the EPR pairs can happen either sequentially or in parallel, depending on the characteristics of the underlying quantum technology. have been generated, the orchestrator retains one ebit of each EPR pair while distributing the other ebit to the corresponding client. Subsequently, once the EPR ebits have been successfully received by the clients, the multipartite entangled state is teleported at the clients by performing local quantum operations (at the orchestrator, plus some correction unitaries at the clients after having received the classical bits) and classical communications (LOCC) [13], as exemplified in Fig. 2. In the following, we assume the LOCC involved in the teleporting protocol to be noiseless. Under this hypothesis, it has been proved that the quantum teleportation is deterministic, as long as the entangled pairs distributed between the orchestrator and clients are maximally entangled [27]. In other words, if EPR pairs are successfully generated and distributed between orchestrator and clients, the multipartite state itself is consequently successfully distributed via quantum teleportation. Clearly, entanglement purification strategies can be proactively exploited for achieving maximally entangled pairs. The rational for assuming noiseless LOCC lies in the focus of the manuscript, which aims at analysing how the transduction impacts multipartite entanglement distribution. Consequently, since in TMD strategies the transduction process acts only on the EPR pairs while the multipartite state to be distributed remains “untouched", we neglect the noise in the LOCC that are not directly involved in the transduction problem [^5] We refer the reader to [13], for a detailed analysis about the impact of noisy LOCC on quantum teleportation in presence of superconducting nodes. In the aforementioned work, it has been showed, by exploiting the IBM quantum platform [37], that noisy LOCC affects the fidelity of the teleported state..
Accordingly, TMD shifts the impact of noisy quantum transduction and noisy optical-ebit carrier propagation from multipartite ebits to the ebits of the EPRs. Indeed, no up- or down-conversion of the multipartite ebits is required, since the multipartite state is involved only in local operations for implementing the quantum teleportation process [27, 38, 39], while the ebits to be transduced are those of the EPR pairs. Consequently, i) if one of the ebits is lost (corrupted) due to the noisy transduction and/or noisy fiber transmission, only the corresponding EPR pair needs to be regenerated, rather than the entire multipartite state. This is much more easy; ii) this shifting makes TMD viable for all the classes of multipartite entanglement, regardless of their persistence properties.
In nutshell, TMD strategy guarantees more resilience to noise and better protection against decoherence [4, 40], since any loss or noise would affect the EPR pair only.